OpenStudy (logan13):

What is the area of the square above? 6 inches2 13 inches2 15 inches2 9 inches2

OpenStudy (logan13):

https://app159.studyisland.com/images/blank.gif

OpenStudy (unanimoose):

You are just so sad.

OpenStudy (anonymous):

Sides?

OpenStudy (logan13):

@AlexandervonHumboldt2

OpenStudy (anonymous):

Tell me sides?

OpenStudy (alexandervonhumboldt2):

area=side^2

OpenStudy (unanimoose):

It's a square so just do \[a^2\] and you have the answer.

OpenStudy (unanimoose):

Just multiply one side by itself and you have the area.

OpenStudy (anonymous):

All sides of a square are equal So let's say that one side is equal to 4. Then the area would be: 4 x 4 = 16 so 16 inches ^2 but for your example i dont have sides so im gonna go with D, 9 inches ^2

OpenStudy (unanimoose):

Now please try and do the rest yourself instead of posting every question on here.

OpenStudy (logan13):

3in 3in

OpenStudy (alexandervonhumboldt2):

ok so 3^2=???

OpenStudy (unanimoose):

\[3^2 = ?\]

OpenStudy (anonymous):

You would do 3 x 3 = 9 because remember all sides of a square are equal

OpenStudy (anonymous):

D, 9 inches ^2